# Sparse factorization of the square all-ones matrix of arbitrary order

### v1

@@ 5 Sequential doubly stochastic factorization
$A_{\mathrm{L}}$ $:=\widehat{T}^{(1)}\widehat{T}^{(2)}\cdots\widehat{T}^{(\tau)}$ (23a)
@@
$A_{\mathrm{R}}$ $:=\widehat{T}^{(\tau)}\widehat{T}^{(\tau-1)}\cdots\widehat{T}^{(1)}$ (23b)
$\;=(I_{n_{1}}\oplus\cdots\oplus(I_{n_{\tau-1}}\oplus(I_{n_{\tau}}\oplus T^{(\tau)})\cdot T^{(\tau-1)})\cdot T^{(\tau-2)})\cdots T^{(1)}.$ (23c)
@@
$\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}V^{(k)}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}=\frac{1}{m_{k-1}}\mathds{1}_{m_{k-1}}\mathds{1}_{m_{k-1}}^{\mathsf{T}},$ (25)
@@
$=\begin{bmatrix}J_{\tau-1}&\\ &J_{\tau}\end{bmatrix}T^{(\tau-1)}\begin{bmatrix}J_{\tau-1}&\\ &J_{\tau}\end{bmatrix}$ (26j)
$=\begin{bmatrix}J_{\tau-1}&\\ &J_{\tau}\end{bmatrix}\begin{bmatrix}\widetilde{A}^{(\tau-1)}_{11}&\widetilde{A}^{(\tau-1)}_{12}\\ \big{(}\widetilde{A}^{(\tau-1)}_{12}\big{)}^{\mathsf{T}}&\tfrac{m_{\tau-1}}{m_{\tau-2}}I\end{bmatrix}\begin{bmatrix}J_{\tau-1}&\\ &J_{\tau}\end{bmatrix}$ (26q)
$=\begin{bmatrix}J_{\tau-1}\widetilde{A}^{(\tau-1)}_{11}J_{\tau-1}&J_{\tau-1}\widetilde{A}^{(\tau-1)}_{12}J_{\tau}\\ \big{(}J_{\tau-1}\widetilde{A}^{(\tau-1)}_{12}J_{\tau}\big{)}^{\mathsf{T}}&\tfrac{m_{\tau-1}}{m_{\tau-2}}J_{\tau}^{2}\end{bmatrix}$ (26t)
$=\frac{1}{m_{\tau-2}}\mathds{1}_{m_{\tau-2}}\mathds{1}_{m_{\tau-2}}^{\mathsf{T}}.$ (26u)
The first equation (26e) uses $\mkern 1.5mu\overline{\mkern-1.5muJ}_{\tau-2}=J_{\tau-1}\oplus\mkern 1.5mu\overline{\mkern-1.5muJ}_{\tau-1}=J_{\tau-1}\oplus J_{\tau}$. Then, (26j) and (26q) use the definition $T^{(\tau)}=I_{n_{\tau}}$ and (21). Finally, (26u) follows from the updates (16) and (17) in Algorithm 2, as well as the fact $J_{\tau}^{2}=J_{\tau}$.
@@
$=\begin{bmatrix}J_{k-1}&\\ &\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}\end{bmatrix}\begin{bmatrix}\widetilde{A}^{(k-1)}_{11}&\widetilde{A}^{(k-1)}_{12}\\ \big{(}\widetilde{A}^{(k-1)}\big{)}^{\mathsf{T}}&\tfrac{m_{k-1}}{m_{k-2}}I\end{bmatrix}\begin{bmatrix}I_{n_{k-1}}&\\ &V^{(k)}\end{bmatrix}\begin{bmatrix}J_{k-1}&\\ &\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}\end{bmatrix}$ (27i)
$=\begin{bmatrix}J_{k-1}A^{(k-1)}_{11}J_{k-1}&J_{k-1}A^{(k-1)}_{12}V^{(k)}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}\\ \big{(}J_{k-1}A^{(k-1)}_{12}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}\big{)}^{\mathsf{T}}&\tfrac{m_{k-1}}{m_{k-2}}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}V^{(k)}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}\end{bmatrix}$ (27l)
$=\begin{bmatrix}\tfrac{1}{m_{k-2}}\mathds{1}_{n_{k-1}}\mathds{1}_{n_{k-1}}^{\mathsf{T}}&\tfrac{1}{m_{k-2}}\mathds{1}_{n_{k-1}}\mathds{1}_{m_{k-1}}^{\mathsf{T}}\\ \tfrac{1}{m_{k-2}}\mathds{1}_{m_{k-1}}\mathds{1}_{n_{k-1}}^{\mathsf{T}}&\tfrac{1}{m_{k-2}}\mathds{1}_{m_{k-1}}\mathds{1}_{m_{k-1}}^{\mathsf{T}}\end{bmatrix}$ (27o)
$=\frac{1}{m_{k-2}}\mathds{1}_{m_{k-2}}\mathds{1}_{m_{k-2}}^{\mathsf{T}}.$
In (27i), we use $\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-2}=J_{k-1}\oplus\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}$, the definition of $T^{(k)}$ in (21), and the definition of direct sum. After multiplying out all the matrices in (27l), the third equation (27o) follows from the definition of $A^{(k-1)}_{11}$ and $A^{(k-1)}_{12}$ (see (16)–(17)), the definition of $V^{(k)}$ in (24), and the assumption that identity (25) holds for $k\in[\tau-1]$. In particular, the $(1,2)$-subblock of (27l) is further simplified as follows:
@@
$=\frac{1}{n_{k-1}}\mathds{1}_{n_{k-1}}\left(\mathds{1}_{n_{k-1}}^{\mathsf{T}}\begin{bmatrix}\operatorname{diag}\Big{(}\tfrac{n_{k-1}}{m_{k-2}}\mathds{1}_{m_{k-1}}\Big{)}\\ 0\end{bmatrix}\right)V^{(k)}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}$ (28c)
$=\frac{1}{m_{k-2}}\mathds{1}_{n_{k-1}}\big{(}\mathds{1}_{m_{k-1}}^{\mathsf{T}}V^{(k)}\big{)}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}$ (28d)
$=\frac{1}{m_{k-2}}\mathds{1}_{n_{k-1}}\mathds{1}_{m_{k-1}}^{\mathsf{T}}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}$ (28e)
$=\frac{1}{m_{k-2}}\mathds{1}_{n_{k-1}}\mathds{1}_{m_{k-1}}^{\mathsf{T}}.$ (28f)
In (28c), we use the definition of $\widetilde{A}^{(k-1)}_{12}$ in (17), and (28d) writes out $\mathds{1}_{n_{k-1}}^{\mathsf{T}}\widetilde{A}^{(k-1)}_{12}=\frac{n_{k-1}}{m_{k-2}}\mathds{1}_{m_{k-1}}^{\mathsf{T}}$. Then, (28e) and (28f) use the doubly stochastic property of $V^{(k)}$ and $\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}$, respectively.
@@
The first equation uses the convention $\mkern 1.5mu\overline{\mkern-1.5muJ}_{0}=J_{0}$ and the relation $A_{\mathrm{L}}=T^{(1)}(I_{n_{1}}\oplus V^{(2)})=V^{(1)}$. The second equation applies (25) with $k=1$, and the last one follows from the convention $m_{0}=n$. Then, the right SDS factorization follows directly from the fact that $A_{\mathrm{R}}=A_{\mathrm{L}}^{\mathsf{T}}$ and thus
$J=(J_{0}A_{\mathrm{L}}J_{0})^{\mathsf{T}}=J_{0}A_{\mathrm{L}}^{\mathsf{T}}J_{0}=J_{0}A_{\mathrm{R}}J_{0}.$
Finally, the doubly stochastic property of $A_{\mathrm{L}}$ (and $A_{\mathrm{R}}$) follows from that of $\{T^{(k)}\}$ and the fact that the product of doubly stochastic matrices is still doubly stochastic. ∎
@@
$d_{\mathrm{max}}(A_{\mathrm{R}})$ $=d_{\mathrm{max}}((V^{(1)})^{\mathsf{T}})=2d_{\mathrm{max}}(V^{(2)})=\cdots=2^{\tau-1}d_{\mathrm{max}}(V^{(\tau)})=2^{\tau-1}.$
6 Application in decentralized average consensus and optimization
@@ 6.1 Decentralized average consensus
We first formulate the decentralized average consensus problem as follows. In a group of $n$ agents, each one holds a piece of information, denoted by $x^{(0)}_{i}\in\mathbb{R}^{d}$, and the entire group aims to compute the average $\mkern 1.5mu\overline{\mkern-1.5mux}:=\tfrac{1}{n}\sum_{i=1}^{n}x^{(0)}_{i}$ via communication. The communication (or connection) between agents is modeled by a sequence of (undirected) graphs (or topologies) ${\mathcal{G}}^{(k)}=({\mathcal{V}},W^{(k)},{\mathcal{E}}^{(k)})$, where ${\mathcal{V}}=\{1,\ldots,n\}$ is the vertex set representing agents and each ${\mathcal{E}}^{(k)}\subseteq{\mathcal{V}}\times{\mathcal{V}}$ is the set of edges (or connections). It is assumed that the set of agents remains static while the set of edges can be time-varying. The entry $w_{ij}^{(k)}\in\mathbb{R}_{\geq 0}$ in the mixing matrix $W^{(k)}$ applies a weighting factor to the information exchanged between agent $j$ and agent $i$. If $w_{ij}^{(k)}=0$, it means agent $i$ is not a neighbor of agent $j$ in ${\mathcal{G}}^{(k)}$; i.e., $(i,j)\notin{\mathcal{E}}^{(k)}$. The state of agent $i$ (or the information held by agent $i$) at iteration $k$ is designated as $x_{i}^{(k)}$ and evolves according to the following recursion: for $k\in\mathbb{N}$,
@@
    The limit of each $x_{i}^{(k)}$ is $\mkern 1.5mu\overline{\mkern-1.5mux}$:
    $\lim_{k\to\infty}x^{(k)}_{i}=\mkern 1.5mu\overline{\mkern-1.5mux},\ \ \text{for all}\ i\in[n].$
@@
In modern application scenarios involving GPUs and high-performance computing (HPC) resources, we can design and alter the communication topology in the decentralized averaging process. In this case, sparse factorization of $J$ helps design topologies with cheap communication cost per iteration and achieve consensus within a finite number of iterations (29). To see this, consider a set of sparse matrices $\{W^{(i)}\}^{q}_{i=1}$ that satisfies $J=W^{(q)}\cdots W^{(2)}W^{(1)}$. When the associated graph sequence $\{{\mathcal{G}}^{(i)}\}_{i=1}^{q}$ is used as the (time-varying) topologies for decentralized averaging, the iteration (29) yields
@@
Therefore, unlike classical results where consensus is achieved only asymptotically, finite-time consensus (i.e., consensus in exactly $q$ iterations) in decentralized averaging is achieved by exploiting sparse factorization of $J$.
To this end, the proposed HB and SDS factorizations of $J$ can be used to construct sparse graph sequences with the desirable finite-time consensus property for arbitrary number of agents $n\in\mathbb{N}_{\geq 2}$. This is in contrast to most previous work, which has requirements on the matrix order $n$ (e.g., $n=p^{\tau}$ for some $(p,\tau)\in\mathbb{N}_{\geq 2}\times\mathbb{N}_{\geq 1}$). Below, we describe in detail how to exploit the factorization $J=J_{0}AJ_{0}$ to construct graph sequences $\{W^{(i)}\}$ with finite-time consensus, and then discuss two additional advantages of the proposed HB and SDS factorizations.
@@
In addition to the ability to handle an arbitrary number of agents, the proposed factorizations (RHB, DSHB, SDS) provide more flexibility to balance the communication costs and the number of communication rounds toward consensus. Recall that the communication cost involved in the decentralized averaging iteration (29) is related to the total number of nonzeros and the largest node degree in the communication graph. Thus, using sparser mixing matrices would reduce communication costs but likely increase the total number of averaging iterations toward consensus. For example, using the left (or right) SDS factor $A_{\mathrm{L}}$ (or $A_{\mathrm{R}}$) completes Phase 2 in one iteration, while using the “one-peer” $T$-factors in (21) results in $\tau-1$ iterations in Phase 2. Such a trade-off in the choice of Phase-2 matrices is summarized in Table 1.
Table 1: Trade-offs between the communication cost (modeled by the largest node degree $d_{\mathrm{max}}$) and the number of iterations in phase 2.
@@
Num. of nonzeros $n+\tau(\tau-1)$ $\sum\limits_{k=1}^{\tau}kn_{k}$ $\sum\limits_{k=1}^{\tau}(2^{k}-1)n_{k}$ $\sum\limits_{k=1}^{\tau}(2^{k}-1)n_{k}$ $n_{k}+2\sum\limits_{i=k+1}^{\tau}n_{i}$
@@
Moreover, the proposed form of factorization (2) can resolve a practical concern omitted by classical settings of decentralized averaging (and optimization). It is typically assumed that the distance between agents is equidistant and that each agent is indistinguishable from another. In practice, however, it may not be the case. Consider the scenario where agents are HPC resources. Agents A and B may be allocated on the same physical machine while agent C is on another physical machine. Consequently, the communication cost between agents A and B is cheaper than that between agents A and C. It is natural to model the network as several sub-networks where each sub-network is a cluster of relatively “close” agents. Such a structure is easily exploited by factorization $J=J_{0}AJ_{0}$. Communication in phases 1 and 3 is all intra-cluster and can be modeled by different sparse factorizations of $J_{k}$, $k\in[\tau]$. The more expensive inter-cluster communication only happens in Phase 2 and is modeled by the sparse matrix $A$, which can be the RHB factor, the DSHB factor, the SDS factors ($A_{\mathrm{L}}$ or $A_{\mathrm{R}}$), or a sequence of $T$-factors in (21). The proposed factorization form (2) promotes cheap, intra-cluster communications and limits the more expensive, inter-cluster ones.

### v2

@@ 5 Sequential doubly stochastic factorization
$A_{\mathrm{L}}$ $:=\widehat{T}^{(1)}\widehat{T}^{(2)}\cdots\widehat{T}^{(\tau)}$
@@
$A_{\mathrm{R}}$ $:=\widehat{T}^{(\tau)}\widehat{T}^{(\tau-1)}\cdots\widehat{T}^{(1)}$
$\;=(I_{n_{1}}\oplus\cdots\oplus(I_{n_{\tau-1}}\oplus(I_{n_{\tau}}\oplus T^{(\tau)})\cdot T^{(\tau-1)})\cdot T^{(\tau-2)})\cdots T^{(1)}.$
@@
$\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}V^{(k)}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}=\tfrac{1}{m_{k-1}}\mathds{1}_{m_{k-1}}\mathds{1}_{m_{k-1}}^{\mathsf{T}},$ (24)
@@
$=\begin{bmatrix}J_{\tau-1}&\\ &J_{\tau}\end{bmatrix}T^{(\tau-1)}\begin{bmatrix}J_{\tau-1}&\\ &J_{\tau}\end{bmatrix}$ (25b)
$=\begin{bmatrix}J_{\tau-1}&\\ &J_{\tau}\end{bmatrix}\begin{bmatrix}\widetilde{A}^{(\tau-1)}_{11}&\widetilde{A}^{(\tau-1)}_{12}\\ \big{(}\widetilde{A}^{(\tau-1)}_{12}\big{)}^{\mathsf{T}}&\tfrac{m_{\tau-1}}{m_{\tau-2}}I\end{bmatrix}\begin{bmatrix}J_{\tau-1}&\\ &J_{\tau}\end{bmatrix}$ (25c)
$=\begin{bmatrix}J_{\tau-1}\widetilde{A}^{(\tau-1)}_{11}J_{\tau-1}&J_{\tau-1}\widetilde{A}^{(\tau-1)}_{12}J_{\tau}\\ \big{(}J_{\tau-1}\widetilde{A}^{(\tau-1)}_{12}J_{\tau}\big{)}^{\mathsf{T}}&\tfrac{m_{\tau-1}}{m_{\tau-2}}J_{\tau}^{2}\end{bmatrix}$
$=\tfrac{1}{m_{\tau-2}}\mathds{1}_{m_{\tau-2}}\mathds{1}_{m_{\tau-2}}^{\mathsf{T}}.$ (25d)
The first equation (25a) uses $\mkern 1.5mu\overline{\mkern-1.5muJ}_{\tau-2}=J_{\tau-1}\oplus\mkern 1.5mu\overline{\mkern-1.5muJ}_{\tau-1}=J_{\tau-1}\oplus J_{\tau}$. Then, (25b) and (25c) use the definition $T^{(\tau)}=I_{n_{\tau}}$ and (21). Finally, (25d) follows from the updates (16) and (17) in Algorithm 2, as well as the fact $J_{\tau}^{2}=J_{\tau}$.
@@
$=\begin{bmatrix}J_{k-1}&\\ &\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}\end{bmatrix}\begin{bmatrix}\widetilde{A}^{(k-1)}_{11}&\widetilde{A}^{(k-1)}_{12}\\ \big{(}\widetilde{A}^{(k-1)}\big{)}^{\mathsf{T}}&\tfrac{m_{k-1}}{m_{k-2}}I\end{bmatrix}\begin{bmatrix}I_{n_{k-1}}&\\ &V^{(k)}\end{bmatrix}\begin{bmatrix}J_{k-1}&\\ &\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}\end{bmatrix}$ (26a)
$=\begin{bmatrix}J_{k-1}A^{(k-1)}_{11}J_{k-1}&J_{k-1}A^{(k-1)}_{12}V^{(k)}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}\\ \big{(}J_{k-1}A^{(k-1)}_{12}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}\big{)}^{\mathsf{T}}&\tfrac{m_{k-1}}{m_{k-2}}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}V^{(k)}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}\end{bmatrix}$ (26b)
$=\begin{bmatrix}\tfrac{1}{m_{k-2}}\mathds{1}_{n_{k-1}}\mathds{1}_{n_{k-1}}^{\mathsf{T}}&\tfrac{1}{m_{k-2}}\mathds{1}_{n_{k-1}}\mathds{1}_{m_{k-1}}^{\mathsf{T}}\\ \tfrac{1}{m_{k-2}}\mathds{1}_{m_{k-1}}\mathds{1}_{n_{k-1}}^{\mathsf{T}}&\tfrac{1}{m_{k-2}}\mathds{1}_{m_{k-1}}\mathds{1}_{m_{k-1}}^{\mathsf{T}}\end{bmatrix}$ (26c)
$=\tfrac{1}{m_{k-2}}\mathds{1}_{m_{k-2}}\mathds{1}_{m_{k-2}}^{\mathsf{T}}.$
In (26a), we use $\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-2}=J_{k-1}\oplus\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}$, the definition of $T^{(k)}$ in (21), and the definition of direct sum. After multiplying out all the matrices in (26b), the third equation (26c) follows from the definition of $A^{(k-1)}_{11}$ and $A^{(k-1)}_{12}$ (see (16)–(17)), the definition of $V^{(k)}$ in (23), and the assumption that identity (24) holds for $k\in[\tau-1]$. In particular, the $(1,2)$-block of (26b) is further simplified as follows:
@@
$=\tfrac{1}{n_{k-1}}\mathds{1}_{n_{k-1}}\left(\mathds{1}_{n_{k-1}}^{\mathsf{T}}\begin{bmatrix}\operatorname{diag}\Big{(}\tfrac{n_{k-1}}{m_{k-2}}\mathds{1}_{m_{k-1}}\Big{)}\\ 0\end{bmatrix}\right)V^{(k)}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}$ (27a)
$=\tfrac{1}{m_{k-2}}\mathds{1}_{n_{k-1}}\big{(}\mathds{1}_{m_{k-1}}^{\mathsf{T}}V^{(k)}\big{)}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}$ (27b)
$=\tfrac{1}{m_{k-2}}\mathds{1}_{n_{k-1}}\mathds{1}_{m_{k-1}}^{\mathsf{T}}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}$ (27c)
$=\tfrac{1}{m_{k-2}}\mathds{1}_{n_{k-1}}\mathds{1}_{m_{k-1}}^{\mathsf{T}}.$ (27d)
In (27a), we use the definition of $\widetilde{A}^{(k-1)}_{12}$ in (17), and (27b) writes out $\mathds{1}_{n_{k-1}}^{\mathsf{T}}\widetilde{A}^{(k-1)}_{12}=\frac{n_{k-1}}{m_{k-2}}\mathds{1}_{m_{k-1}}^{\mathsf{T}}$. Then, (27c) and (27d) use the doubly stochastic property of $V^{(k)}$ and $\mkern 1.5mu\overline{\mkern-1.5muJ}_{k-1}$, respectively.
@@
The first equation uses the convention $\mkern 1.5mu\overline{\mkern-1.5muJ}_{0}=J_{0}$ and the relation $A_{\mathrm{L}}=T^{(1)}(I_{n_{1}}\oplus V^{(2)})=V^{(1)}$. The second equation applies (24) with $k=1$, and the last one follows from the convention $m_{0}=n$. Then, the right SDS factorization follows directly from the fact that $A_{\mathrm{R}}=A_{\mathrm{L}}^{\mathsf{T}}$ and thus $J=(J_{0}A_{\mathrm{L}}J_{0})^{\mathsf{T}}=J_{0}A_{\mathrm{L}}^{\mathsf{T}}J_{0}=J_{0}A_{\mathrm{R}}J_{0}$. Finally, the doubly stochastic property of $A_{\mathrm{L}}$ (and $A_{\mathrm{R}}$) follows from that of $\{T^{(k)}\}$ and the fact that the product of doubly stochastic matrices is still doubly stochastic. ∎
@@
$d_{\mathrm{max}}(A_{\mathrm{R}})$ $=d_{\mathrm{max}}((V^{(1)})^{\mathsf{T}})=2d_{\mathrm{max}}(V^{(2)})=\cdots=2^{\tau-1}d_{\mathrm{max}}(V^{(\tau)})=2^{\tau-1}.\qed$
6 Application in decentralized averaging and optimization
@@
The application scenario considered here involves abstracting agents as high-performance computing (HPC) resources. In modern data centers, machines are organized (or clustered) into racks, and switches and routers are used to connect machines physically. These connections form the so-called physical topology, which is often fixed [16, §2]. In comparison, virtual topology refers to the logical network layout between virtual machines (VMs) and other HPC resources. It models how data are transferred between VMs and has the following properties (see, e.g., [1]).
  • Flexibility. Virtual topology can be dynamically redesigned without reconfiguring the physical hardware.
  • Scalability. Virtual topology can span multiple physical data centers.
  • Automation. Management of virtual topology can be automated via software-defined network (SDN) controllers.
These properties are crucial for adapting to evolving traffic demands with the purpose of, e.g., reducing power consumption, network congestion, end-to-end delay, or blocking probability. Hence, with virtual topology, the connection between machines (agents) can be easily manipulated in a dynamic manner. However, the communication cost between machines is based on their physical locations. For example, communication within a network segment (e.g., a rack or a section of the data center) is often more efficient and economical than that between network segments [16, §2]. The proposed sparse factorization (2) of $J$ offers a robust solution for designing dynamic virtual topologies as it exploits the clustering and hierarchical structure in data centers, leverages the flexibility of virtual topology, and takes into account the non-uniform communication costs within a virtual topology.
@@ 6.1 Decentralized average consensus
The decentralized average consensus (or decentralized averaging) problem can be formally formulated as follows. In a group of $n$ agents, each one holds a piece of information, denoted by $x^{(0)}_{i}\in\mathbb{R}^{d}$, and the entire group aims to compute the average $\mkern 1.5mu\overline{\mkern-1.5mux}:=\tfrac{1}{n}\sum_{i=1}^{n}x^{(0)}_{i}$ via communication. The communication (or connection) between agents is modeled by a sequence of (undirected) graphs (or topologies) $\mathcal{G}^{(k)}=(\mathcal{V},W^{(k)},\mathcal{E}^{(k)})$, where $\mathcal{V}=\{1,\ldots,n\}$ is the node set representing agents, each $\mathcal{E}^{(k)}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges (or connections), and the weighted adjacency matrix $W^{(k)}\in\mathbb{R}^{n\times n}$ stores the weights of the edges. It is assumed that the set of agents remains static while the set of edges can be time-varying. In this context, $W^{(k)}$ is often called the mixing matrix, and its entry $w_{ij}^{(k)}\in\mathbb{R}_{\geq 0}$ applies a weighting factor to the information exchanged between agent $j$ and agent $i$. If $w_{ij}^{(k)}=0$, it means agent $i$ is not a neighbor of agent $j$ in $\mathcal{G}^{(k)}$; i.e., $(i,j)\notin\mathcal{E}^{(k)}$. We do not distinguish between a graph $\mathcal{G}$ and its weight matrix $W$. The state of agent $i$ (or the information held by $i$) at iteration $k$ is designated as $x_{i}^{(k)}$ and evolves according to the following recursion: for $k\in\mathbb{N}$,
@@
    The limit of each $x_{i}^{(k)}$ is $\mkern 1.5mu\overline{\mkern-1.5mux}$: $\lim_{k\to\infty}x^{(k)}_{i}=\mkern 1.5mu\overline{\mkern-1.5mux}$ for all $i\in[n]$.
@@
In this section, we are interested in modern applications where virtual machines in HPC scenarios are abstracted as agents. In this case, the communication topology can be customized and easily altered during the averaging process. Hence, the proposed sparse factorization of $J$ helps design topologies that achieve consensus within a finite number of communication rounds. To see this, consider a set of sparse matrices $\{W^{(i)}\}^{q}_{i=1}$ that satisfies (1). When the associated graph sequence $\{\mathcal{G}^{(i)}\}_{i=1}^{q}$ is used as the (time-varying) topologies for decentralized averaging, the iteration (28) yields
@@
Therefore, unlike classical results where consensus is achieved only asymptotically, finite-time consensus (i.e., consensus in exactly $q$ communication rounds) in decentralized averaging is achieved by exploiting sparse factorization of $J$. Moreover, the sparsity of all the factors in the proposed factorizations of $J$ helps reduce the per-round communication costs. In decentralized average consensus (as well as decentralized optimization), the communication cost at each round is modeled by either the total number of nonzeros in the mixing matrix $W$ or the largest node degree $d_{\mathrm{max}}$; see, e.g., the recent book [13, §11.3].
To this end, the proposed HB and SDS factorizations of $J$ can be used to construct sparse graph sequences with cheap per-round communication costs and the desirable finite-time consensus property for arbitrary number of agents $n\in\mathbb{N}_{\geq 2}$. This is in contrast to most previous work, which has requirements on the matrix order $n$ (e.g., $n=p^{\tau}$ for some $(p,\tau)\in\mathbb{N}_{\geq 2}\times\mathbb{N}_{\geq 1}$). Below, we describe in detail how to exploit the factorization $J=J_{0}AJ_{0}$ to construct graph sequences $\{W^{(i)}\}$ with finite-time consensus, and then discuss two additional advantages of the proposed HB and SDS factorizations.
@@
In addition to the ability to handle an arbitrary number of agents, the proposed factorizations (RHB, DSHB, SDS) provide more flexibility to balance the communication costs and the number of communication rounds toward consensus. Recall that the communication cost involved in each iteration (28) is related to the total number of nonzeros and the largest node degree in the communication topology. Thus, using sparser graphs would reduce communication costs but likely increase the total number of iterations toward consensus. For example, using $A_{\mathrm{L}}$ (or $A_{\mathrm{R}}$) completes Phase 2 in one iteration, while using the “one-peer” $T$-factors in (21) results in $\tau-1$ iterations in Phase 2. Such a trade-off is summarized in Table 1.
@@
Num. of nonzeros $n+\tau(\tau-1)$ $\sum_{k=1}^{\tau}kn_{k}$ $\sum_{k=1}^{\tau}(2^{k}-1)n_{k}$ $\sum_{k=1}^{\tau}(2^{k}-1)n_{k}$ $n_{k}+2\sum_{i=k+1}^{\tau}n_{i}$
@@
Table 1: Trade-offs between the communication cost (modeled by either the number of nonzeros or the largest node degree $d_{\mathrm{max}}$) and the number of communication rounds in Phase 2.
Moreover, the proposed form of factorization (2) handles the issue of non-uniform communication costs mentioned at the beginning of Section 6. In classical decentralized settings, it is typically assumed that the distance between agents is equidistant and that each agent is indistinguishable from another. However, this is not the case in the virtual topology of modern data centers. Recall that inter-cluster communication between machines in a rack (or a section of the data center) is often cheaper and swifter than intra-cluster communication. Such a characteristic is easily exploited by factorization $J=J_{0}AJ_{0}$. Communication in Phases 1 and 3 is all intra-cluster and can be modeled by different sparse factorizations of $J_{k}$, $k\in[\tau]$. The more expensive inter-cluster communication only happens in Phase 2 and is modeled by the sparse matrix $A$. So, the proposed factorization form (2) promotes cheap, intra-cluster communications and limits the more expensive, inter-cluster ones.
Numerical experiments
Here, we verify that the proposed sparse factors of $J$ satisfy the finite-time consensus property (1) in numerical experiments. To do so, we simulate an average consensus problem. Each agent is initialized with a random vector $x_{i}^{(0)}\sim\mathcal{N}(0,\Sigma)$ drawn from a Gaussian distribution (with $\Sigma$ symmetric positive definite). The iterates $x_{i}^{(k)}$ evolve according to the recursion (28), and the consensus error at each iteration is defined as $\Xi^{(k)}:=\tfrac{1}{n}\sum_{i=1}^{n}\|x_{i}^{(k)}-\mkern 1.5mu\overline{\mkern-1.5mux}\|^{2}$.
Figure 2 presents the consensus error using the proposed graph sequences, the one-peer exponential graphs [21] and $p$-peer hyper-cuboids [12] for various numbers of agents. (The experimental settings and the presentation style strictly follow from [21, 12].) It is known that when $n$ is not a power of $2$, using the one-peer exponential graphs cannot achieve finite-time consensus [21]. This result, together with Theorems 1, 2 and 3, is verified numerically in Figure 2. More specifically, the proposed factorizations have a steep drop in the consensus error, indicating the vanishing of the consensus error, while for one-peer exponential graphs, the consensus error decreases asymptotically. Moreover, note that the number of communication rounds is not the only criterion considered in practice. The per-round communication cost is also important. For example, when $n=43$ (a prime number), the $p$-peer hyper-cuboid reduces to the fully-connected graph and reaches consensus in just one round, while SDS takes 12 rounds. However, the total communication cost (measured by total number of nonzeros; see Table 1) for $p$-peer hyper-cuboid is $1849$, compared to only 488 for the SDS factorization.
Figure 2: Consensus error versus the number of communication rounds. We examine five graph sequences over four different numbers of agents. The results for SDS factorization, the one-peer exponential graphs, and the $p$-peer hyper-cuboids are presented on the left; the results for RHB and DSHB factorizations are presented on the right. For SDS, RHB, and DSHB, the partition of $n$ (3) and the factors used in each cluster are listed at the right-most of legend. For example, $[[2,2,2,2,2],[2,2,2],[2]]$ means the 42 agents are partitioned into three clusters with sizes 32, 8, and 2 and each cluster is further binary partitioned for $J_{0}$ generation. The factors used for p-peer hyper-cuboid are $42=[2,3,7]$, 43=[43]43delimited-[]4343=[43]43 = [ 43 ], $254=[2,127]$, and $255=[3,5,17]$ respectively.